OpenStudy (sonyalee77):

Please help!!!! 4x^2+6/2 *3/2x^2+3

OpenStudy (sonyalee77):

@campbell_st

OpenStudy (sonyalee77):

@Savannah_Noelle

OpenStudy (sonyalee77):

@Ashleyisakitty @Euler271

OpenStudy (hihi67):

Okay, I might be able to help you:D

OpenStudy (sonyalee77):

thanx

OpenStudy (savannah_noelle):

sorry I'm not good with math :/ @agent0smith

OpenStudy (campbell_st):

well if you factor the 1st quadratic you'll get \[\frac{2(2x^2 +3)}{2} \times \frac{3}{(2x^2 + 3)}\] you should be able to identify some common factors... that cancel to 1